INFRARED SPECTROSCOPY OF W(CO)_{6} IN CRYOGENIC SOLID PARAHYDROGEN

\underline{GARIMELLA V. SUBRAHMANYAM} and DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

We are currently engaged in studying the infrared spectroscopy of W(CO)_{6} in cryogenic solid parahydrogen in the CO stretch region. This work builds off the beautiful infrared spectroscopic and photochemical studies of W(CO)_{6} conducted in rare gas matrices. Infrared spectra in the CO stretch region were measured and analyzed in order to study how the vibrational modes of the W(CO)_{6} are perturbed in the low temperature solid hydrogen environment. In this regard we will present the infrared spectral data and discuss our current understanding of this spectrum.